Solve by the elimination method
-4x-5y=2
12x+20y=-9
and
5x-y=24
x+4y=9

Answer by Alan3354(69443)

Solve by the elimination method
-4x-5y=2
12x+20y=-9
Multiply the 1st eqn by 3, then add.
====================
and
5x-y=24
x+4y=9
Multiply the 1st eqn by 4, then add.